If the base of a triangle is increased by 20% and its height is doubled, by what percent will its area be increased?
Answer by stanbon(75887)

If the base of a triangle is increased by 20% and its height is doubled, by what percent will its area be increased?
-----
Original Area = (1/2)(base)(height)
-----
New Area = (1/2)(1.2*base)(2h) = (1/2)(1.2)(2)(base)(height)
----
= 2.4[(1/2)(base)(height) = (1/2)(bh) + 1.2(1/2)(bh)
-----
Percent increase = 120%
-----------------------------
Cheers,
Sta H.